A small company has 20 employees. Six of these employees will be selected randomly to be interviewed as part of an employee satisfaction program. How many different groups of six can be selected?



1
Expert's answer
2021-06-08T14:23:30-0400
"\dbinom{20}{6}=\dfrac{20!}{6!(20-6)!}=\dfrac{20(19)(18)(17)(16)(15)}{1(2)(3)(4)(5)(6)}=38760"